OpenStudy (anonymous):

Algebra Review Question:

OpenStudy (anonymous):

\[2\sqrt{3}\div \sqrt{6}\]

OpenStudy (jdoe0001):

\(\bf 2\sqrt{3}\div \sqrt{6}\implies \cfrac{2\sqrt{3}}{\sqrt{6}}\implies \cfrac{2\sqrt{3}}{\sqrt{2\cdot 3}}\implies \cfrac{2\cancel{ \sqrt{3} }}{\sqrt{2}\cancel{ \sqrt{3} }}\)

OpenStudy (jdoe0001):

and then you could simplify that further if you wish

OpenStudy (anonymous):

Final answer would be \[\frac{ 2\sqrt{2} }{ 2 }\] right?

OpenStudy (jdoe0001):

yeap \(\bf 2\sqrt{3}\div \sqrt{6}\implies \cfrac{2\sqrt{3}}{\sqrt{6}}\implies \cfrac{2\sqrt{3}}{\sqrt{2\cdot 3}}\implies \cfrac{2\cancel{ \sqrt{3} }}{\sqrt{2}\cancel{ \sqrt{3} }} \\ \quad \\ \cfrac{2}{\sqrt{2}}\implies \cfrac{\cancel{ 2 }\sqrt{2}}{\cancel{ 2 }}\)